DragonSlaYeR:

Evaluate the expression -3m + 5 when m = 2 Evaluate the expression 4x - 3y when x = 10 and y = -2 I know this has been a challenging school year so far for many students and families. I see so many of you working hard to attend our Math Google Meets, complete the assignments, ask for help when you have questions, and retake assignments when needed. Thank you for your efforts! I also know many of you have family members or other people who have helped support you with your math learning and I want to send a special thank you message to them. Please answer the following questions to help me determine who to thank: Do you do most of your math work on your own or do you have someone at home who helps you understand our lessons better? If someone helps you, what is their name and how they are related to you? If someone helps you, how often or how many hours do you think you work together in a week?

DragonSlaYeR:

@dontsaymyname

DragonSlaYeR:

plz help

jhonyy9:

Evaluate the expression -3m + 5 when m = 2 Evaluate the expression 4x - 3y when x = 10 and y = -2 so do you know what mean evaluate these expressions ?

DragonSlaYeR:

naw

DragonSlaYeR:

there r three questions

DragonSlaYeR:

...

DragonSlaYeR:

help me plz

dontsaymyname:

-3(2) +5 = what? 4(10) - 3( -2) = what?

jhonyy9:

how you evaluate -3m +5 when m = 2 ?

DragonSlaYeR:

-1

jhonyy9:

yes

DragonSlaYeR:

and then 46

DragonSlaYeR:

wb I know this has been a challenging school year so far for many students and families. I see so many of you working hard to attend our Math Google Meets, complete the assignments, ask for help when you have questions, and retake assignments when needed. Thank you for your efforts! I also know many of you have family members or other people who have helped support you with your math learning and I want to send a special thank you message to them. Please answer the following questions to help me determine who to thank: Do you do most of your math work on your own or do you have someone at home who helps you understand our lessons better? If someone helps you, what is their name and how they are related to you? If someone helps you, how often or how many hours do you think you work together in a week?

dontsaymyname:

That you're going to have to answer on your own because I dont know who helps you with your math

DragonSlaYeR:

no one lol

DragonSlaYeR:

my parents don't know this stuff

dontsaymyname:

Then say you do most of your school work on your own :)

DragonSlaYeR:

lol